# Unifying Simulation and Inference with Normalizing Flows

### v1

@@ I Introduction
Over the last years, some partial solutions to these two central challenges have been proposed. Generalized numerical inversion [26, 27, 28] is a prior-independent calibration approach, although it does not scale well to many output dimensions. The Gaussian Ansatz [17, 18] is a local maximum likelihood estimator, and thus prior-independent (more in Sec. II). For UQ, loss functions can be modified to estimate quantiles in addition to just the mean, median or mode [13, 14] and deep generative models can estimate the exact resolution function [24, 25]. The Gaussian Ansatz produces a local estimate of the resolution and so is currently the only method that is both prior-independent and that goes beyond a point estimate. However, the Gaussian Ansatz only gives a Gaussian approximation of the resolution. We note that it is also possible to use Bayesian versions [29, 30, 31, 32, 33] of neural networks to go beyond point estimates by providing the uncertainty of the output. However, the uncertainty would still be prior-dependent.
In this paper, we propose a prior-independent method for detector calibrations that produces per-shower resolution estimates. The approach is based on deep generative models with access to the explicit likelihood. In particular, we use normalizing flows [^2] It is also possible to extract the likelihood in many dimensions from a diffusion model [34] or if the inputs are discretized, from an autoregressive model [35]. [36, 37, 38, 39, 40, 41]. These machine learning models are invertible functions with a tractable Jacobian so that one can be used for both sampling and density estimation. The core idea behind a normalizing flow is that one starts with a simple random variable with a known probability density (e.g. a Gaussian) and then applies a series of transformations with the tractable Jacobian. This results in a complex probability density that can also be computed via the change of variables formula. The probability density can be made conditional by letting each transformation depend on the conditional quantity.
@@
    With access to the complete likelihood, we can compute the exact resolution function. This allows us to obtain per-shower resolution estimates which give us additional information about how close the predicted energy is to the true energy. Furthermore, it enables us to estimate asymmetries in the resolution function (not possible with the GA).
@@ II Methods
    The bias of a calibration is the deviation between the central tendency of the inferred estimate $\hat{z}$ and the true reference value $z$. Any measure of central tendency can be used to measure closure, such as the median or mode. In this paper, we will focus on the mode [^3] Usually bias refers to the mean, but experimentalists typically calibrate to the mode. of the inferred estimates given a fixed $z$ value. To compute the mode in practice, for a given $z$, we look at $K$ measured values $x_{1},x_{2}\dots x_{K}$ corresponding to that $z$. Each measured value $x_{i}$ undergoes calibration to derive an inferred estimate $\hat{z}_{i}$. The mode is subsequently determined based on the collection of inferred estimates $\{\hat{z}_{1},\hat{z}_{2},\dots,\hat{z}_{K}\}$.
@@
    • Full resolution: For fixed $z$, it is defined as half of the 68% confidence interval about the mode of a collection of inferred estimates.
@@ II.2 Maximum Likelihood Inference
Also, the challenge with MLE is that we usually do not know $p_{X|Z}$ explicitly. Nevertheless, we are able to sample from this conditional density by running a simulation. The Gaussian Ansatz [17, 18] approaches this problem by estimating the likelihood locally around the mode with a clever neural network parameterization and learning rate schedule. In this work, we instead aim to learn the entire likelihood using a neural network — an approach sometimes referred to as neural likelihood estimation in the simulation-based inference literature [78, 79, 80, 81]. This is substantially more work, but we observe that the work may already be done: given a fast simulation based on neural networks with access to the likelihood, we can use it for calibration in addition to generation without requiring any additional retraining.
@@ III.4 Resolution estimation
MLE-based calibration methods, such as the NF and Gaussian Ansatz, are able to provide estimates of the per-shower resolution and this allows these methods to go beyond a point estimate by providing per-event resolutions (i.e. per-shower resolutions in calorimeter context). This gives MLE-based calibration an important advantage over direct regression.
@@
Since the NF is able to access the complete likelihood, this enables it to capture non-Gaussian components of the resolution. This is one advantage it has over the the Gaussian Ansatz approach  [17, 18] — the only other method that is both prior-independent and can provide a per-shower resolution estimate — which is, by construction, unable to predict any asymmetry. We quantify this by defining an asymmetry observable based on the upper and lower bounds of the 68% confidence interval of the MLE prediction. In particular, the asymmetry is computed as the ratio $\frac{\Delta_{L}}{\Delta_{U}}$, where $\Delta_{L}$ ($\Delta_{U}$) is the difference between the lower (upper) bound of the 68% confidence interval and the MLE.
An example is included in Fig. 6 which shows the per-shower resolution asymmetry predicted by the NF for 100k showers with $E_{\rm true}=50$ GeV. We observe that $\frac{\Delta_{L}}{\Delta_{U}}$ is largely greater than unity. This implies that for most showers, lower $E_{\rm inc}$ values tend to have larger likelihoods. While one might suspect that this resolution asymmetry is merely an artifact of the flow, we observe in Fig. 7 that the distribution of $E_{\rm inc}$ at fixed vertical slices of $E_{\rm dep}$ is also asymmetric towards lower values of $E_{\rm inc}$. The particular shape of the 2D histogram in Fig. 7 is due to the nature of the $\pi^{+}$ showers. At fixed $E_{\rm inc}$, fully electromagnetic events will have larger $E_{\rm dep}$, since all the energy is deposited via electromagnetic interactions. On the other hand, fully hadronic events will have significantly lower $E_{\rm dep}$ due to energy carried away invisibly by particles like neutrons and neutrinos produced in nuclear interactions. The $\pi^{+}$ showers considered in this work contain both electromagnetic and hadronic components, so on a shower-by-shower basis, the $E_{\rm dep}$ can range anywhere between these two extremes of a fully electromagnetic shower and a fully hadronic shower [91]. This results in a tail at high energies in the $E_{\rm dep}$ distribution for fixed $E_{\rm inc}$, presumably because the typical $\pi^{+}$ shower is mostly hadronic in nature. Thinking of $p(E_{\rm dep}|E_{\rm inc})$ as a proxy for the likelihood suggests that the resolution asymmetry is not merely an artifact of the flow but a genuine characteristic to be expected. We hypothesize that the ability to predict the exact resolution would allow the NF to outperform the Gaussian Ansatz in resolution estimation in calibration tasks where the asymmetries in the resolution are important.
@@ IV Conclusions
We demonstrate a second advantage of using a MLE-based calibration methods over direct regression by estimating per-shower resolutions. The average estimated per-shower resolution obtained from CaloFlow closely aligns with the full resolution. In contrast, the direct regression approach yields only point estimates and lacks the capacity to make per-shower resolution predictions. For CaloFlow, we found that smaller per-shower resolutions tend to coincide with more accurately calibrated showers. This is evidence that the per-shower resolutions from CaloFlow are reliable and have the potential to be leveraged for improved calibration. Notably, utilizing a NF for calibration grants access to the complete resolution function, a feature inaccessible in previous methods like the Gaussian Ansatz. This characteristic enables the NF to capture asymmetries in the resolution, thereby expected to confer an advantage over the Gaussian Ansatz in specific calibration tasks. In future work, it would be interesting to study how the per-shower resolution information obtained by the NF can be utilized to further improve the calibration.
@@ Acknowledgements
We would like to thank Rikab Gambhir for the helpful discussions related to the Gaussian Ansatz model. CK would like to thank the Baden-Württemberg-Stiftung for financing through the program Internationale Spitzenforschung, project Uncertainties – Teaching AI its Limits (BWST_IF2020-010). IP and DS are supported by the U.S. Department of Energy (DOE), Office of Science grant DOE-SC0010008 and HD, VM, and BN are supported by the DOE under contract DE-AC02-05CH11231.

### v2

@@ I Introduction
Over the last years, some solutions to these two central challenges have been proposed. Generalized numerical inversion [26, 27, 28] is a prior-independent calibration approach, although it does not scale well to many output dimensions. The Gaussian Ansatz [17, 18] is a maximum likelihood estimator, and thus prior-independent (more in Sec. II). For UQ, loss functions can be modified to estimate quantiles in addition to just the mean, median or mode [13, 14] and deep generative models can estimate the exact resolution function [24, 25]. The Gaussian Ansatz produces an estimate of the resolution and so is currently the only method that is both prior-independent and that goes beyond a point estimate. We note that it is also possible to use Bayesian versions [29, 30, 31, 32, 33] of neural networks to go beyond point estimates by providing the uncertainty of the output. However, the uncertainty would still be prior-dependent.
In this paper, we propose a prior-independent method for detector calibrations that produces per-shower resolution estimates. The approach is based on deep generative models with access to the explicit likelihood. In particular, we use normalizing flows [^2] It is also possible to extract the likelihood in many dimensions from a diffusion model [34] or if the inputs are discretized, from an autoregressive model [35]. [36, 37, 38, 39, 40, 41]. These machine learning models are invertible functions with a tractable Jacobian so that one can be used for both sampling and density estimation. The core idea behind a normalizing flow is that one starts with a simple random variable with a known probability density (e.g. a Gaussian) and then applies a series of transformations with the tractable Jacobian. This results in a complex probability density that can also be computed via the change of variables formula. The probability density can be made conditional by letting each transformation depend on the conditional quantity. Our approach serves as an alternative to other methods based on maximum likelihood estimation.
@@
    With access to the complete likelihood, we can compute the exact resolution function. This allows us to obtain per-shower resolution estimates which give us additional information about how close the predicted energy is to the true energy. Furthermore, it enables us to estimate asymmetries in the resolution function.
@@ II Methods
    The bias of a calibration is the deviation between the central tendency of the inferred estimate $\hat{z}$ and the true reference value $z$. Any measure of central tendency can be used to measure closure, such as the median or mode. In this paper, we will focus on the mode [^3] Usually bias refers to the mean, but experimentalists typically calibrate to the mode. of the inferred estimates given a fixed $z$ value. To compute the mode in practice, for a given $z$, we look at $K$ measured values $x_{1},x_{2}\dots x_{K}$ corresponding to that $z$. Each measured value $x_{i}$ undergoes calibration to derive an inferred estimate $\hat{z}_{i}$. The mode is subsequently determined based on the collection of inferred estimates $\{\hat{z}_{1},\hat{z}_{2},\dots,\hat{z}_{K}\}$ (see steps in Sec. III.3).
@@
    • Full resolution: For fixed $z$, it is defined as half of the 68% confidence interval about the mode determined based on a collection of inferred estimates.
@@ II.2 Maximum Likelihood Inference
Also, the challenge with MLE is that we usually do not know $p_{X|Z}$ explicitly. Nevertheless, we are able to sample from this conditional density by running a simulation. In this work, we instead aim to learn the entire likelihood using a neural network — an approach sometimes referred to as neural likelihood estimation in the simulation-based inference literature [78, 79, 80, 81]. This is substantially more work, but we observe that the work may already be done: given a fast simulation based on neural networks with access to the likelihood, we can use it for calibration in addition to generation without requiring any additional retraining.
@@ III.4 Resolution estimation
MLE-based calibration methods, such as the NF and Gaussian Ansatz, are able to provide estimates of the per-shower resolution and this allows these methods to go beyond a point estimate by providing per-event resolutions (i.e. per-shower resolutions in calorimeter context). This gives MLE-based calibration an important advantage over direct regression. Unlike the Gaussian Ansatz, the NF serves as a generative model, presenting a notable advantage over the Gaussian Ansatz: a NF utilized for surrogate modeling can subsequently be repurposed for calibration tasks without additional retraining.
@@
Since the NF is able to access the complete likelihood, this enables it to capture non-Gaussian [^8] In principle, the Gaussian Ansatz [17, 18] should also be able to capture non-Gaussian components of the resolution. However, this has yet to be shown in a worked example. Previous works [18, 91] only attempted to compute the Gaussian resolution. Also, Gaussian Ansatz is trained in a different way (see Ref. [17]) compared to the NF and is more sensitive to hyperparameter choices. components of the resolution. We quantify this by defining an asymmetry observable based on the upper and lower bounds of the 68% confidence interval of the MLE prediction. In particular, the asymmetry is computed as the ratio $\frac{\Delta_{L}}{\Delta_{U}}$, where $\Delta_{L}$ ($\Delta_{U}$) is the difference between the lower (upper) bound of the 68% confidence interval and the MLE.
An example is included in Fig. 6 which shows the per-shower resolution asymmetry predicted by the NF for 100k showers with $E_{\rm true}=50$ GeV. We observe that $\frac{\Delta_{L}}{\Delta_{U}}$ is largely greater than unity. This implies that for most showers, lower $E_{\rm inc}$ values tend to have larger likelihoods. While one might suspect that this resolution asymmetry is merely an artifact of the flow, we observe in Fig. 7 that the distribution of $E_{\rm inc}$ at fixed vertical slices of $E_{\rm dep}$ is also asymmetric towards lower values of $E_{\rm inc}$. The particular shape of the 2D histogram in Fig. 7 is due to the nature of the $\pi^{+}$ showers. At fixed $E_{\rm inc}$, fully electromagnetic events will have larger $E_{\rm dep}$, since all the energy is deposited via electromagnetic interactions. On the other hand, fully hadronic events will have significantly lower $E_{\rm dep}$ due to energy carried away invisibly by particles like neutrons and neutrinos produced in nuclear interactions. The $\pi^{+}$ showers considered in this work contain both electromagnetic and hadronic components, so on a shower-by-shower basis, the $E_{\rm dep}$ can range anywhere between these two extremes of a fully electromagnetic shower and a fully hadronic shower [92]. This results in a tail at high energies in the $E_{\rm dep}$ distribution for fixed $E_{\rm inc}$, presumably because the typical $\pi^{+}$ shower is mostly hadronic in nature. Thinking of $p(E_{\rm dep}|E_{\rm inc})$ as a proxy for the likelihood suggests that the resolution asymmetry is not merely an artifact of the flow but a genuine characteristic to be expected.
@@ IV Conclusions
We demonstrate a second advantage of using a MLE-based calibration methods over direct regression by estimating per-shower resolutions. The average estimated per-shower resolution obtained from CaloFlow closely aligns with the full resolution. In contrast, the direct regression approach yields only point estimates and lacks the capacity to make per-shower resolution predictions. For CaloFlow, we found that smaller per-shower resolutions tend to coincide with more accurately calibrated showers. This is evidence that the per-shower resolutions from CaloFlow are reliable and have the potential to be leveraged for improved calibration. Notably, utilizing a NF for calibration grants access to the complete resolution function. This characteristic enables the NF to capture asymmetries in the resolution. In future work, it would be interesting to study how the per-shower resolution information obtained by the NF can be utilized to further improve the calibration.
@@ Acknowledgements
We would like to thank Rikab Gambhir and Jesse Thaler for the helpful discussions related to the Gaussian Ansatz model and for feedback on the manuscript. CK would like to thank the Baden-Württemberg-Stiftung for financing through the program Internationale Spitzenforschung, project Uncertainties – Teaching AI its Limits (BWST_IF2020-010). IP and DS are supported by the U.S. Department of Energy (DOE), Office of Science grant DOE-SC0010008 and HD, VM, and BN are supported by the DOE under contract DE-AC02-05CH11231.